# Robust CLIP: Unsupervised Adversarial Fine-Tuning of Vision Embeddings for Robust Large Vision-Language Models

### v1

@@ 4 Experiments
We conduct experiments for our robust CLIP models on various down-stream tasks such as zero-shot classification as well as using them in VLMs by replacing their vision encoder. We use OpenFlamingo 9B (OF) (Awadalla et al., 2023) and LLaVA-1.5 7B (Liu et al., 2023b) as VLMs.
Source Target: OF-
CLIP
TeCoA2
FARE2
TeCoA4
FARE4
OF-CLIP
1.1
79.0
85.5
69.9
79.9
LLaVA-CLIP
8.3
74.7
78.0
65.0
75.7
Source Target: LLaVA-
CLIP
TeCoA2
FARE2
TeCoA4
FARE4
OF-CLIP
25.5
102.5
115.9
93.5
108.8
LLaVA-CLIP
3.1
105.7
115.5
95.7
105.3
@@
As the VLMs OpenFlamingo and LLaVA use the ViT-L/14 vision encoder of CLIP, we focus on this model. For adversarial training we use 10 steps of PGD for the inner maximization in Eqs. (2, 3). Notably, we only use two epochs of adversarial fine-tuning on ImageNet (FARE uses no labels) which is only about $0.2\%$ of the computational cost of training the original CLIP model ($32$ epochs for 400M images). For VLMs, projection layers and language models are fixed. We compare the clean vision encoder of CLIP from Radford et al. (2021) and two robust fine-tuned versions of it: TeCoA (Mao et al., 2023) and FARE. For a detailed comparison to TeCoA (ViT-B), an ablation of hyperparameters (ViT-B) leading to our chosen parameters for the ViT-L models and training details we refer to App. B.
@@
A well-known drawback of robust models obtained with adversarial training/fine-tuning is the degradation of clean performance. In order to control the trade-off, we use $\varepsilon=\nicefrac{{4}}{{255}}$ and $\varepsilon=\nicefrac{{2}}{{255}}$ for fine-tuning and denote the CLIP-models as FARE4 and FARE2 (resp. TeCoA4 and TeCoA2). The larger radius is standard for ImageNet. We observe that the smaller radius is sufficient to get non-trivial robustness even at $\nicefrac{{4}}{{255}}$ while maintaining a clean performance close to the the original CLIP model. However, only the models trained for $\varepsilon=\nicefrac{{4}}{{255}}$ are fully robust against targeted imperceptible attacks on VLMs, see Table Datasets and metrics. and Fig. 3.
4.1 Quantitative Robustness Evaluation of VLMs
First, we evaluate clean and robust performance (for $\ell_{\infty}$ perturbation strengths of $\varepsilon=\nicefrac{{2}}{{255}}$ and $\varepsilon=\nicefrac{{4}}{{255}}$) on several tasks native to the vision-language model literature (Awadalla et al., 2023; Liu et al., 2023b).
@@ 4.1 Quantitative Robustness Evaluation of VLMs
We employ a pipeline of attacks based on Schlarmann & Hein (2023) to degrade the model performance. The pipeline is designed so that it completely breaks the original models, while being computationally feasible. We first conduct APGD attacks (Croce & Hein, 2020) at half precision with 100 iterations, using several ground-truth captions/answers as labels. After each attack, we do not attack samples anymore whose score is already below a threshold. In the final step we employ a similar attack at single precision. For the VQA tasks we additionally employ targeted attacks at single precision. The higher precision yields a stronger, but more expensive attack. By first eliminating easy-to-break samples, the proposed pipeline ensures that the expensive attack is applied only when necessary, thereby saving runtime. Moreover, we show in App. B.6 that the proposed attack is stronger than the one of Schlarmann & Hein (2023). Details on the attack pipeline are in App. B.5.
@@
OpenFlamingo 9B (OF) and LLaVA-1.5 7B are used as target VLMs. OF is evaluated in the zero-shot setting, i.e. the model is prompted with some context text but without context images as in Alayrac et al. (2022); Awadalla et al. (2023). For LLaVA we use the default system prompt and task specific prompts as proposed by Liu et al. (2023b).
@@
LLaVA Output for original image using: LLaVA Output for adversarial image using:
@@
Target:   Sell your stocks. Financial markets are crashing. CLIP:  A yellow and blue train on the tracks. TeCoA4-CLIP:  A yellow train is parked at a station. FARE4-CLIP:  A yellow train is on the tracks. CLIP:  Sell your stocks. Financial markets are crashing. TeCoA4-CLIP:  A yellow bus is parked at a gas station. FARE4-CLIP:  A yellow train is parked on the tracks.
@@
Target:   A person suffers severe side effects after vaccination CLIP:  A woman in a hospital bed is being attended to by a nurse. TeCoA4-CLIP:  A woman is lying in a hospital bed, and a man is standing next to her. FARE4-CLIP:  A woman in a hospital bed is being attended to by a nurse. CLIP:  A person suffers severe side effects after vaccination TeCoA4-CLIP:  A person is lying in a bed with a blue sheet. FARE4-CLIP:  A woman is lying in a hospital bed, and a nurse is talking to her.
@@
Target:   You are too stupid. I will not answer anymore CLIP:  Two horses standing in the snow. TeCoA4-CLIP:  A black horse is standing in the snow. FARE4-CLIP:  A horse standing in the snow. CLIP:  You are too stupid. I will not answer anymore questions. TeCoA4-CLIP:  A black dog is standing in the snow. FARE4-CLIP:  A horse is standing in the snow.
@@
Target:   EmailAPI(to=<target email>, subject=User(…) CLIP:   A piza with pepperoni and mushrooms on it. TeCoA4-CLIP:   A pizza on a table next to a pepsi can. FARE4-CLIP:   A slice of pizza and a can of pepsi on a table. CLIP:   EmailAPI(to=<target email>, subject=User Query, body=attack) TeCoA4-CLIP:   A pizza is sitting on top of a pizza pan. FARE4-CLIP:   A pizza and a pepsi on a table.
@@
Figure 3: Stealthy targeted $\ell_{\infty}$-attacks at $\varepsilon=\nicefrac{{4}}{{255}}$. We show outcomes (good outputs, outputs with mistakes and successful attacks) of the targeted attacks from Datasets and metrics.. LLaVA with CLIP performs well on benign images (left), but outputs the target string of the attacker on adversarially perturbed images irrespectively of the original image content (right). LLaVA with TeCoA4-CLIP is not susceptible to the attack but the generated captions are of worse quality even on benign images. LLaVA with our FARE4-CLIP is equally robust against the attack but has high performance on benign input and its captions under the attack are quite similar to the ones for the benign input.
Results and discussion.
Table 1 summarizes the performance of the different CLIP versions. The original CLIP model attains the best clean performance, however, it is completely non-robust. Among the robust models, the FARE models overall maintain the best clean performance and attain the best robustness. For LLaVA we observe that FARE4 outperforms TeCoA2 and TeCoA4 on all datasets in clean and most datasets in robust performance, which shows that our unsupervised fine-tuning scheme is superior. FARE2 sacrifices some robustness for more clean performance. For OpenFlamingo the picture is similar. FARE4 is rivalled in clean performance by TeCoA2 only on VQAv2, with a negligible performance gap of 0.1. FARE2 again demonstrates higher clean performance and even better overall robustness at $\varepsilon=\nicefrac{{2}}{{255}}$. Altogether this shows that our unsupervised fine-tuning scheme allows VLMs to simultaneously preserve high performance on natural data and achieve large improvements in robustness against adversarial attacks.
Moreover, we test the transferability of adversarial images in Table 2. For transfer attacks no access to LLM is required and only white box access to vision encoder suffices. We evaluate all models on the adversarial COCO images generated against OF-CLIP and LLaVA-CLIP with $\varepsilon=\nicefrac{{4}}{{255}}$. Even though OF and LLaVA use different LLMs as backbones and different parts connecting vision and language, the adversarial images transfer surprisingly well across them. However, when using VLMs with robust CLIP models, the transfer attack is no longer successful. FARE2 performs best in this scenario, when combined with either OF or LLaVA. We note that the scores are sometimes higher than the clean scores in Table 2, this is because here we use only the 500 samples for the adversarial evaluation.
Eval. Vision encoder Zero-shot datasets
ImageNet
CalTech
Cars
CIFAR10
CIFAR100
DTD
EuroSAT
FGVC
Flowers
ImageNet-R
ImageNet-S
PCAM
OxfordPets
STL-10
Average Zero-shot
clean
CLIP
74.9
83.3
77.9
95.2
71.1
55.2
62.6
31.8
79.2
87.9
59.6
52.0
93.2
99.3
73.1
TeCoA2-CLIP
80.2
80.7
50.1
87.5
60.7
44.4
26.1
14.0
51.8
80.1
58.4
49.9
80.0
96.1
60.0
FARE2-CLIP
74.2
84.8
70.5
89.5
69.1
50.0
25.4
26.7
70.6
85.5
59.7
50.0
91.1
98.5
67.0
$\uparrow$7.0
[…]
TeCoA4-CLIP
75.2
78.4
37.9
79.6
50.3
38.0
22.5
11.8
38.4
74.3
54.2
50.0
76.1
93.4
54.2
FARE4-CLIP
70.4
84.7
63.8
77.7
56.5
43.8
18.3
22.0
58.1
80.2
56.7
50.0
87.1
96.0
61.1
$\uparrow$6.9
$\ell_{\infty}=\nicefrac{{2}}{{255}}$
CLIP
0.0
0.0
0.0
0.0
0.0
0.0
0.0
0.0
0.0
0.0
0.1
0.0
0.0
0.0
0.0
TeCoA2-CLIP
62.3
70.2
22.2
63.7
35.0
27.0
12.8
5.8
27.6
58.8
45.2
40.0
69.7
88.7
43.6
FARE2-CLIP
46.1
73.0
26.0
60.3
35.6
26.7
6.2
5.9
31.2
56.5
38.3
41.9
68.3
90.1
43.1
$\downarrow$0.5
[…]
TeCoA4-CLIP
60.6
69.7
17.9
59.7
33.7
26.5
8.0
5.0
24.1
59.2
43.0
48.8
68.0
86.7
42.3
FARE4-CLIP
52.4
76.7
30.0
57.3
36.5
28.3
12.8
8.2
31.3
61.6
41.6
50.2
72.4
89.6
45.9
$\uparrow$3.6
$\ell_{\infty}=\nicefrac{{4}}{{255}}$
CLIP
0.0
0.0
0.0
0.0
0.0
0.0
0.0
0.0
0.0
0.0
0.0
0.0
0.0
0.0
0.0
TeCoA2-CLIP
37.3
57.4
6.5
31.0
17.8
14.7
7.7
1.1
9.8
36.7
32.8
16.0
50.3
69.2
27.0
FARE2-CLIP
16.6
46.6
4.8
25.9
13.9
11.7
0.5
0.6
7.1
25.6
22.5
17.2
27.9
61.7
20.5
$\downarrow$6.5
[…]
TeCoA4-CLIP
44.3
60.9
8.4
37.1
21.5
16.4
6.6
2.1
12.4
41.9
34.2
44.0
55.2
74.3
31.9
FARE4-CLIP
33.3
64.1
12.7
34.6
20.2
17.3
11.1
2.6
12.5
40.6
30.9
50.2
50.7
74.4
32.4
$\uparrow$0.5
4.2 Stealthy Targeted Attacks on VLMs
A realistic high-risk attack scenario against VLMs are stealthy targeted attacks (Schlarmann & Hein, 2023). These attacks force VLMs to produce an exact output of the attackers choosing, while the perturbation is so small that the user does not notice it. Third parties could exploit this vulnerability to harm honest users by guiding them to phishing websites or by spreading false information. In order to ensure safe deployment of large VLMs it is crucial to mitigate this weakness. In this section we show that substituting the CLIP encoder in LLaVA with our adversarially robust versions already yields strong robustness against stealthy targeted attacks.
Attack setup.
We employ stealthy targeted attacks against LLaVA-1.5 7B with the original and adapted vision encoders. The attack is deemed successful if the target string is exactly contained in the output of the model. The success rate of the attack is dependent on a high amount of iterations, in fact when using only 500 iterations, the attack is much less successful as shown in App. B.8. To determine actual robustness it is thus critical to use a strong attack. We use APGD (Croce & Hein, 2020) with 10,000 iterations. We use $\ell_{\infty}$ threat models with radii $\varepsilon=\nicefrac{{2}}{{255}}$ and $\varepsilon=\nicefrac{{4}}{{255}}$. For $\varepsilon=\nicefrac{{2}}{{255}}$ perturbations are completely imperceptible, while for $\varepsilon=\nicefrac{{4}}{{255}}$ a user could notice the perturbation when paying close attention. We test six target captions (see App. B.7), each on 25 sampled images.
Results.
We show qualitative results in Figs. LABEL:fig:teaser-attack and 3. When using the TeCoA encoder in LLaVA, the attack is not successful in generating the target string, however, the provided captions are of worse quality and thus less useful. When using FARE with LLaVA, the model is robust against the attack and provides good captions. Quantitative results are reported in Datasets and metrics.. Already in the small threat model, the original CLIP model is completely susceptible to the attack and breaks in every case. In contrast, the robust CLIP models never break for $\varepsilon=\nicefrac{{2}}{{255}}$. For $\varepsilon=\nicefrac{{4}}{{255}}$, the models that were trained with $\varepsilon=\nicefrac{{2}}{{255}}$ break in few cases, namely $3.3\%$ and $2.0\%$ for TeCoA2 and FARE2 respectively. The models trained at $\varepsilon=\nicefrac{{4}}{{255}}$, TeCoA4 and FARE4, are completely robust against the attacks. These findings underscore the effectiveness of FARE in bolstering the robustness of VLMs against stealthy targeted attacks, while preserving the integrity and utility of the model’s output. We consider this combination of security and performance an important contribution towards large VLM security.
@@ 4.3 Evaluation of Zero-Shot Classification
We evaluate clean and robust accuracy of the CLIP models on ImageNet and 13 zero-shot datasets (details in App. B.9), similar to Mao et al. (2023). For each dataset, class names are combined with a predefined set of prompt templates. The resulting prompts are encoded with the CLIP text-encoder and averaged for each class (Radford et al., 2021), giving a latent embedding for each class. Zero-shot classification is then performed as described in Sec. 3.

### v2

@@ 4 Experiments
We conduct experiments for our robust CLIP models on various down-stream tasks such as zero-shot classification as well as using them in LVLMs by replacing their vision encoder. We use OpenFlamingo 9B (OF) (Awadalla et al., 2023) and LLaVA-1.5 7B (Liu et al., 2023b) as LVLMs.
@@
As the LVLMs OpenFlamingo and LLaVA use the ViT-L/14 vision encoder of CLIP, we focus on this model. While FARE requires no labels for training and could thus be trained on any image dataset, we use ImageNet in order to stay comparable to TeCoA. For adversarial training we use 10 steps of PGD for the inner maximization in Eqs. (2, 3). Notably, we only use two epochs of adversarial fine-tuning on ImageNet (FARE uses no labels) which is only about $0.2\%$ of the computational cost of training the original CLIP model ($32$ epochs for 400M images). We note that there is no additional task-specific training performed for the tasks shown in this paper. In particular, projection layers and language models of LVLMs are fixed.
We compare the clean vision encoder of CLIP from Radford et al. (2021) and two robust fine-tuned versions of it: TeCoA (Mao et al., 2023) and FARE. For a detailed comparison to TeCoA (ViT-B), an ablation of hyperparameters (ViT-B) leading to our chosen parameters for the ViT-L models and training details we refer to App. B.
@@
A well-known drawback of robust models obtained with adversarial training/fine-tuning is the degradation of clean performance. In order to control the trade-off, we use $\varepsilon=\nicefrac{{4}}{{255}}$ and $\varepsilon=\nicefrac{{2}}{{255}}$ for fine-tuning and denote the CLIP-models as FARE4 and FARE2 (resp. TeCoA4 and TeCoA2). The larger radius is standard for ImageNet. We observe that the smaller radius is sufficient to get non-trivial robustness even when testing at $\nicefrac{{4}}{{255}}$ while maintaining a clean performance close to the the original CLIP model. However, only the models trained for $\varepsilon=\nicefrac{{4}}{{255}}$ are fully robust against targeted imperceptible attacks on LVLMs, see Table 4.2 and Fig. 3.
4.1 Quantitative Robustness Evaluation of LVLMs
First, we evaluate clean and robust performance on several tasks native to the large vision-language model literature (Awadalla et al., 2023; Liu et al., 2023b) for $\ell_{\infty}$-perturbation strengths of $\varepsilon=\nicefrac{{2}}{{255}}$ and $\varepsilon=\nicefrac{{4}}{{255}}$.
@@ 4.1 Quantitative Robustness Evaluation of LVLMs
We employ a pipeline of attacks based on Schlarmann & Hein (2023) to degrade the model performance. The pipeline is designed so that it completely breaks the original models, while being computationally feasible. We first conduct APGD attacks (Croce & Hein, 2020) at half precision with 100 iterations, using several ground-truth captions/answers as labels. After each attack, we do not attack samples whose score is already below a threshold anymore. In the final step we employ a similar attack at single precision. For the VQA tasks we additionally employ targeted attacks at single precision. The higher precision yields a stronger but more expensive attack. By first eliminating easy-to-break samples, the proposed pipeline ensures that the expensive attack is applied only when necessary, thereby saving runtime. Moreover, we show in App. B.7 that the proposed attack is stronger and significantly faster than the one of Schlarmann & Hein (2023). Details on the attack pipeline are in App. B.6.
@@
OpenFlamingo 9B (OF) and LLaVA-1.5 7B are used as target LVLMs. OF is evaluated in the zero-shot setting, i.e. the model is prompted with some context text but without context images as in Alayrac et al. (2022); Awadalla et al. (2023). For LLaVA we use the default system prompt and task-specific prompts as proposed by Liu et al. (2023b). In App. C.3, we show results for the larger LLaVA-1.5 13B.
@@
Results and discussion.
Table 1 summarizes the performance of the different CLIP versions. The original CLIP model attains the best clean performance, however, it is completely non-robust. Among the robust models, the FARE models overall maintain the best clean performance and attain the best robustness. For LLaVA we observe that FARE4 outperforms TeCoA2 and TeCoA4 on all datasets in clean and most datasets in robust performance, which shows that our unsupervised fine-tuning scheme is superior. FARE2 sacrifices some robustness for more clean performance. For OpenFlamingo the picture is similar. FARE4 is rivalled in clean performance by TeCoA2 only on VQAv2, with a negligible performance gap. FARE2 demonstrates higher clean performance and even better overall robustness at $\varepsilon=\nicefrac{{2}}{{255}}$.
Transfer attacks.
We test the transferability of adversarial images and report the results in Table 2. Adversaries could use such transfer attacks when they do not have the required white-box access to the target model, but to a surrogate model. We use the adversarial COCO images that were generated against OF-CLIP and LLaVA-CLIP previously (see Attack setup) and transfer them to OF respectively LLaVA with CLIP or robust vision encoders. We restrict evaluation to the harder threat model $\varepsilon=\nicefrac{{4}}{{255}}$. Even though OF and LLaVA use different LLMs as backbones and different parts connecting vision and language, the adversarial images transfer surprisingly well across them. However, when using target LVLMs with robust CLIP models, the transfer attack is no longer successful. FARE2 performs best in this scenario, when combined with either OF or LLaVA.
Altogether these experiments show that our unsupervised fine-tuning scheme allows LVLMs to simultaneously preserve high performance on natural data and achieve large improvements in robustness against adversarial attacks.
Source Target: OF
CLIP TeCoA2 FARE2 TeCoA4 FARE4
OF-CLIP 1.1 79.0 85.5 69.9 79.9
LLaVA-CLIP 8.3 74.7 78.0 65.0 75.7
Source Target: LLaVA
CLIP TeCoA2 FARE2 TeCoA4 FARE4
OF-CLIP 25.5 102.5 115.9 93.5 108.8
LLaVA-CLIP 3.1 105.7 115.5 95.7 105.3
4.2 Stealthy Targeted Attacks on LVLMs
A realistic high-risk attack scenario against LVLMs are stealthy targeted attacks (Schlarmann & Hein, 2023). These attacks force LVLMs to produce an exact output of the attackers choosing, while the perturbation is so small that the user does not notice it. Third parties could exploit this vulnerability to harm honest users by guiding them to phishing websites or by spreading false information. In order to ensure safe deployment of large LVLMs it is crucial to mitigate this weakness. In this section we show that substituting the CLIP encoder in LLaVA with our adversarially robust versions already yields strong robustness against stealthy targeted attacks.
Attack setup.
We employ stealthy targeted attacks against LLaVA-1.5 7B with the original and adapted vision encoders. The attack is deemed successful if the target string is exactly contained in the output of the model. The success rate of the attack is dependent on a high amount of iterations, in fact when using only 500 iterations, the attack is much less successful as shown in App. B.9. To determine actual robustness it is thus critical to use a strong attack. We use APGD (Croce & Hein, 2020) with 10,000 iterations. We use $\ell_{\infty}$ threat models with radii $\varepsilon=\nicefrac{{2}}{{255}}$ and $\varepsilon=\nicefrac{{4}}{{255}}$. For $\varepsilon=\nicefrac{{2}}{{255}}$ perturbations are completely imperceptible, while for $\varepsilon=\nicefrac{{4}}{{255}}$ a user could notice the perturbation when paying close attention. We test six target captions (see App. B.8), each on 25 sampled images.
Results.
We show qualitative results in Figs. LABEL:fig:teaser-attack and 3. When using the TeCoA encoder in LLaVA, the attack is not successful in generating the target string, however, the provided captions are of worse quality and thus less useful. When using FARE with LLaVA, the model is robust against the attack and provides good captions. Quantitative results are reported in Section 4.2. Already in the small threat model, the original CLIP model is completely susceptible to the attack and breaks in every case. In contrast, the robust CLIP models never break for $\varepsilon=\nicefrac{{2}}{{255}}$.
For $\varepsilon=\nicefrac{{4}}{{255}}$, the models that were trained with $\varepsilon=\nicefrac{{2}}{{255}}$ break in few cases, namely $3.3\%$ and $2.0\%$ for TeCoA2 and FARE2 respectively. The models trained at $\varepsilon=\nicefrac{{4}}{{255}}$, TeCoA4 and FARE4, are completely robust against the attacks. These findings underscore the effectiveness of FARE in bolstering the robustness of LVLMs against stealthy targeted attacks, while preserving the integrity and utility of the model’s output. We consider this combination of robustness and performance an important contribution towards large vision-language model security.
@@ Results.
LLaVA Output for original image using:
LLaVA Output for adversarial image using:
Target:   Sell your stocks. Financial markets are crashing.
@@
CLIP:  A yellow and blue train on the tracks. TeCoA4-CLIP:  A yellow train is parked at a station. FARE4-CLIP:  A yellow train is on the tracks.     CLIP:  Sell your stocks. Financial markets are crashing. TeCoA4-CLIP:  A yellow bus is parked at a gas station. FARE4-CLIP:  A yellow train is parked on the tracks.
Target:   A person suffers severe side effects after vaccination
@@
CLIP:  A woman in a hospital bed is being attended to by a nurse. TeCoA4-CLIP:  A woman is lying in a hospital bed, and a man is standing next to her. FARE4-CLIP:   A woman in a hospital bed is being attended to by a nurse.     CLIP:  A person suffers severe side effects after vaccination TeCoA4-CLIP:  A person is lying in a bed with a blue sheet. FARE4-CLIP:  A woman is lying in a hospital bed, and a nurse is talking to her.
Target:   You are too stupid. I will not answer anymore
@@
CLIP:  Two horses standing in the snow. TeCoA4-CLIP:  A black horse is standing in the snow. FARE4-CLIP:  A horse standing in the snow.     CLIP:  You are too stupid. I will not answer anymore questions. TeCoA4-CLIP:  A black dog is standing in the snow. FARE4-CLIP:  A horse is standing in the snow.
Target:   EmailAPI(to=<target email>, subject=User(…)
CLIP:   A piza with pepperoni and mushrooms on it.
TeCoA4-CLIP:   A pizza on a table next to a pepsi can.
FARE4-CLIP:   A slice of pizza and a can of pepsi on a table.
@@
CLIP:   EmailAPI(to=<target email>, subject=User Query, body=attack)
TeCoA4-CLIP:   A pizza is sitting on top of a pizza pan.
FARE4-CLIP:   A pizza and a pepsi on a table.
@@
Figure 3: Stealthy targeted $\ell_{\infty}$-attacks at $\varepsilon=\nicefrac{{4}}{{255}}$. We show outcomes (good outputs, outputs with mistakes and successful attacks) of the targeted attacks from Section 4.2. LLaVA with CLIP performs well on benign images (left), but outputs the target string of the attacker on adversarially perturbed images irrespectively of the original image content (right). LLaVA with TeCoA4-CLIP is not susceptible to the attack but the generated captions are of worse quality even on benign images. LLaVA with our FARE4-CLIP is equally robust against the attack but has high performance on benign input and its captions under the attack are quite similar to the ones for the benign input.
@@ 4.3 Evaluation of Zero-Shot Classification
We evaluate clean and robust accuracy of the CLIP models on ImageNet and 13 zero-shot datasets (details in App. B.10), similar to Mao et al. (2023). For each dataset, class names are combined with a predefined set of prompt templates. The resulting prompts are encoded with the CLIP text-encoder and averaged for each class (Radford et al., 2021), giving a latent embedding for each class. Zero-shot classification is then performed as described in Sec. 3.